### v1

# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

Abstract
We study stopping rules for stochastic gradient descent (SGD) for convex optimization from the perspective of anytime-valid confidence sequences. Classical analyses of SGD provide convergence guarantees in expectation or at a fixed horizon, but offer no statistically valid way to assess, at an arbitrary time, how close the current iterate is to the optimum. We develop an anytime-valid, data-dependent upper confidence sequence for the weighted average suboptimality of projected SGD, constructed via nonnegative supermartingales and requiring no smoothness or strong convexity. This confidence sequence yields a simple stopping rule that is provably $\varepsilon$-optimal with probability at least $1-\alpha$ and is almost surely finite under standard stochastic approximation stepsizes. To the best of our knowledge, these are the first rigorous, time-uniform performance guarantees and finite-time $\varepsilon$-optimality certificates for projected SGD with general convex objectives, based solely on observable trajectory quantities.
1 Introduction
Stochastic gradient methods are the workhorses of modern optimization and machine learning. Their appeal lies in a rare combination of simplicity, scalability, and robustness to noisy information. Yet, despite decades of research, a persistent gap remains between theory and practice: during an optimization run, practitioners have no data-dependent, statistically valid indication of how close the current iterate is to optimality.
Classical analyses provide valuable insights into the behavior of stochastic gradient descent (SGD), including convergence in expectation, asymptotic rates, and fixed-horizon high-probability bounds. However, these guarantees typically rely on problem-dependent constants, such as smoothness parameters, strong convexity coefficients, and variance bounds, which are unknown in practice, and they offer no protection under adaptive monitoring or early stopping. Thus, when training a model or tuning a hyperparameter, a user is left with the fundamental but unanswered question:
How close am I to the optimum right now?
In practical settings, such as large-scale training, reinforcement learning, scientific computing, or any computation where time or data is limited, this question is more than academic. Stopping too early risks under-optimization, while stopping too late wastes considerable resources. Yet most stopping rules used in practice are heuristic: they monitor validation curves, impose fixed budgets, or rely on experience rather than statistical guarantees.
In this paper, we revisit this classical challenge from a sequential-inference perspective. The key idea is to view the progress of SGD as a stochastic process that can be monitored through anytime-valid confidence sequences (CSs). A confidence sequence is a sequence of intervals that remains valid uniformly over time, even under continuous monitoring and adaptive stopping. By adapting this framework to stochastic optimization, we obtain fully data-dependent, time-uniform upper bounds on the suboptimality gap, requiring neither knowledge of the optimization horizon nor precise knowledge of problem-dependent parameters. These guarantees immediately yield statistically valid, data-dependent stopping rules that certify $\varepsilon$-optimality under arbitrary adaptive monitoring.
1.1 Problem formulation
To formalize the problem, we consider optimization over a closed and convex feasible region. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a nonempty compact convex set, and let $f:\mathcal{X}\to\mathbb{R}$ be a convex objective function. We denote an arbitrary minimizer of $f$ over $\mathcal{X}$ by
$x^{\star}\in\arg\min_{x\in\mathcal{X}}f(x).$
We study the classical projected variant of stochastic gradient descent, defined recursively by
$x_{t+1}=\Pi_{\mathcal{X}}(x_{t}-\eta_{t}g_{t}),$ (1)
where $\Pi_{\mathcal{X}}$ denotes Euclidean projection onto $\mathcal{X}$ and $\{\eta_{t}\}_{t\geq 1}$ is a predictable stepsize sequence. Formally, we define the filtration $\mathcal{F}_{t}:=\sigma(x_{1},g_{1},\ldots,x_{t},g_{t})$, for $t\geq 1$, and assume that $\eta_{t}$ is $\mathcal{F}_{t-1}$ measurable. Because $x_{1}\in\mathcal{X}$ and each update projects back onto $\mathcal{X}$, we have $x_{t}\in\mathcal{X}$ for all $t\geq 1$. In particular, all iterates lie in a bounded region around $x^{\star}$:
$\|x_{t}-x^{\star}\|\;\leq\;R_{x}:=\textrm{diam}(\mathcal{X})<\infty.$
The stochastic gradients $g_{t}$ are assumed to satisfy the following three conditions.
Assumption 1.1 (Stochastic gradients).
For all $t\geq 1$, the stochastic gradient $g_{t}$ satisfies:
  • (i)
    Unbiasedness: $\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}]=\nabla f(x_{t})$, and
  • (ii)
    Conditional second moment bound: $\mathbb{E}[\|g_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq G^{2}$ almost surely.
  • (iii)
    Conditional sub-Gaussian noise: there exists a constant $\sigma^{2}>0$ such that for $\xi_{t}\coloneqq g_{t}-\nabla f(x_{t})$,
    $\mathbb{E}\!\left[\exp\!\big(\lambda\langle u,\xi_{t}\rangle\big)\,\Big|\,\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\sigma^{2}\|u\|^{2}}{2}\right)\qquad\forall u\in\mathbb{R}^{d},\ \forall\lambda\in\mathbb{R}.$ (2)
Assumption 1.1(i) is the standard unbiasedness condition used in stochastic approximation. Assumption 1.1(ii) requires a uniform bound on the conditional second moment of the stochastic gradients, which is classical in nonasymptotic analyses of SGD and ensures that the quadratic term $\|g_{t}\|^{2}$ remains controlled. Finally, Assumption 1.1(iii) postulates that the gradient noise $\xi_{t}\coloneqq g_{t}-\nabla f(x_{t})$ is conditionally sub-Gaussian given the past. This condition provides a quantitative tail bound on the one-step fluctuations of the stochastic gradients and is natural in time-uniform concentration arguments. Moreover, it is satisfied by a broad class of noise models, including bounded, Gaussian, and many light-tailed stochastic gradients frequently encountered in practice. We operate in the standard stochastic approximation regime:
Assumption 1.2 (Stepsizes).
The predictable stepsizes $\{\eta_{t}\}_{t\geq 1}$ satisfy
$\sum_{t=1}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty,\qquad 0<\eta_{t}<\infty.$ (3)
These conditions ensure that the algorithm continues making progress in early iterations while the cumulative noise remains controlled. It is convenient to introduce two standard cumulative quantities from stochastic approximation,
$S_{t}:=\sum_{s=1}^{t}\eta_{s},\qquad V_{t}:=\sum_{s=1}^{t}\eta_{s}^{2},$
which will appear naturally in our analysis. To measure optimization progress, we focus on the weighted average suboptimality,
$\bar{F}_{t}:=\frac{\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)}{S_{t}}.$ (4)
This statistic is classical in stochastic approximation theory and has a direct optimization interpretation. Indeed, defining the corresponding weighted average iterate
$\bar{x}_{t}:=\frac{1}{S_{t}}\sum_{s=1}^{t}\eta_{s}x_{s},$ (5)
convexity of $f$ implies
$f(\bar{x}_{t})-f(x^{\star})\;\leq\;\bar{F}_{t}.$
Thus $\bar{F}_{t}$ simultaneously controls the performance of the averaged iterate $\bar{x}_{t}$ and serves as the natural target for our anytime-valid analysis.
1.2 Contributions
Our goal is to provide anytime-valid, data-dependent guarantees for the progress of projected stochastic gradient descent. Specifically, given the stochastic gradients $\{g_{t}\}$, the predictable stepsizes $\{\eta_{t}\}$, and the observed iterates $\{x_{t}\}$, we construct an upper confidence sequence for the weighted average suboptimality $\bar{F}_{t}$ in (4), which holds uniformly over all times $t\geq 1$. This confidence sequence depends on trajectory-dependent quantities: the stochastic gradient norms, the stepsizes, and the squared-distance process $\|x_{t}-x^{\star}\|^{2}$. While the latter is not directly observable, it is uniformly bounded by $R_{x}^{2}$ due to compactness of $\mathcal{X}$, which immediately yields a fully implementable bound depending only on the realized gradient norms, the stepsizes, and $R_{x}$. To state our contributions precisely, we recall the formal notion of an anytime-valid upper confidence sequence.
Definition 1.3 (Anytime-valid upper confidence sequence).
Let $\{\bar{F}_{t}\}_{t\geq 1}$ be as in (4). A sequence of random variables $\{U_{t}(\alpha)\}_{t\geq 1}$ is an anytime-valid upper confidence sequence for $\{\bar{F}_{t}\}$ if
$\mathbb{P}\!\left(\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\right)\geq 1-\alpha.$ (6)
These uniform-in-time bounds enable a natural data-dependent stopping rule. Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,1)$, we define the stopping time
$\tau_{\varepsilon}:=\inf\left\{t\geq 1:U_{t}(\alpha)\leq\varepsilon\right\},$ (7)
which returns the first iterate that is certified to be $\varepsilon$-optimal.
Main contributions.
Our main contributions can be summarized as follows.
  • (i)
    Anytime-valid confidence sequence. Under Assumptions 1.1 and 1.2, Theorem 2.2 provides an upper confidence sequence $\{U_{t}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ satisfying
    $\mathbb{P}\!\left(\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\right)\geq 1-\alpha.$
    The bound $U_{t}(\alpha)$ is fully data-dependent and remains valid uniformly over time, including under adaptive monitoring and data-dependent choices of stepsizes.
  • (ii)
    Anytime-valid $\varepsilon$-optimal stopping rule. Fix $\varepsilon>0$ and $\alpha\in(0,1)$. Theorem 3.1 proves that the stopping time $\tau_{\varepsilon}$ defined in (7) is almost surely finite under Assumption 1.2, and the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies
    $\mathbb{P}\!\left(f(\bar{x}_{\tau_{\varepsilon}})-f(x^{\star})\leq\varepsilon\right)\geq 1-\alpha.$
    Thus projected SGD, when equipped with the confidence sequence from (i), becomes an anytime-valid optimization procedure with a rigorous finite-time certificate of $\varepsilon$-optimality.
1.3 Related Work
Conceptually, our work lies at the intersection of two developments: (i) time-uniform inference, which provides a framework for hypothesis tests and confidence intervals that yields valid control over errors uniformly in time, and (ii) stochastic approximation, which delivers nonasymptotic convergence guarantees for stochastic gradient methods in convex optimization. The present paper combines these perspectives by adapting nonnegative supermartingale techniques from time-uniform inference to the projected SGD setting. This yields a fully data-dependent, anytime-valid upper confidence sequence for suboptimality, together with a stopping rule that is provably $\varepsilon$-optimal with high probability and almost surely finite under standard stochastic approximation conditions. In what follows, we position our contributions within the broader literature on stochastic approximation, averaging and weighted iterates, optimization stopping criteria, time-uniform concentration, and confidence-sequence methodology.
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [robbins1951stochastic] established almost sure convergence of iterative schemes under diminishing stepsizes, while [kushner2003stochastic] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [polyak1992acceleration]. This asymptotic perspective was sharpened in [duchi2021asymptotic], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [davis2024asymptotic]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established fixed-confidence bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [nemirovski2009robust, lan2012optimal]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias-variance tradeoffs and finite-time rates for averaged iterates [moulines2011non, bach2013non, dieuleveut2016nonparametric]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [rakhlin2012making, shamir2013stochastic, harvey2019tight, harvey2019simple, jain2021making]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [agarwal2012information]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, developing robust stochastic gradient methods under heavy-tailed gradients via techniques such as clipping, truncation, and median-of-means constructions [gorbunov2020stochastic, cutkosky2021high, davis2021low]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper.
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [ville1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [darling1967confidence], as well as boundary-crossing results for martingales [robbins1970boundary]. Modern developments systematize these principles through mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [howard2020time, howard2021time]; see also the broader literature on self-normalized processes [de2009self]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [ramdas2023game, ramdas2025hypothesis]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [waudby2024anytime, waudby2024estimating]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
Stopping rules and certificates in optimization. Stopping criteria play a central practical role in stochastic optimization, yet they remain poorly understood from a statistical perspective. In practice, stochastic gradient methods are typically terminated using heuristic rules, such as monitoring validation loss, gradient norms, iterate differences, or pre-assigned computational budgets. While these rules are often effective empirically, they lack formal guarantees: they do not provide statistically valid certificates of $\varepsilon$-optimality, and their use under adaptive monitoring is generally incompatible with fixed-horizon analyses. Classical convergence results for SGD (whether asymptotic or nonasymptotic) are typically stated either in expectation or at a predetermined iteration horizon, and therefore do not justify early stopping based on observed trajectories. As a result, there is a gap between the widespread use of early stopping in optimization practice and the lack of rigorous, time-uniform performance certificates. The confidence-sequence framework developed in this paper directly addresses this gap by providing a fully data-dependent stopping rule with provable $\varepsilon$-optimality guarantees that remain valid under arbitrary, adaptive monitoring.
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [rakhlin2012making] establish sharp high-probability convergence rates for SGD under strong convexity, bounded stochastic gradients, and fixed stepsizes, with guarantees stated for a pre-specified iteration horizon. As such, they do not support data-dependent stopping times. More recent work [feng2023anytime] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. In contrast to both works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
1.4 Organization and notation
Organization. Section 2 develops our anytime-valid confidence sequence for the weighted average suboptimality of projected SGD. Section 3 introduces the associated $\varepsilon$-optimal stopping rule and analyzes its stopping-time complexity under standard stepsize schedules. Technical lemmas and auxiliary proofs are collected in Appendix A.
Notation. For a vector $\in\mathbb{R}^{d}, x\|x\|$ denotes the Euclidean norm. For a closed convex set $\mathcal{X}\subset\mathbb{R}^{d}$, $\Pi_{\mathcal{X}}(z)$ is the Euclidean projection of $z$ onto $\mathcal{X}$. We write $\langle x,y\rangle$ for the standard inner product. Given a filtration $\{\mathcal{F}_{t}\}$, a sequence $\{X_{t}\}$ is $\mathcal{F}_{t}$-adapted if $X_{t}$ is $\mathcal{F}_{t}$-measurable, and is predictable if $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Expectations conditional on $\mathcal{F}_{t-1}$ are written $\mathbb{E}[\cdot\mid\mathcal{F}_{t-1}]$. For sequences $\{\eta_{t}\}$ and $\{x_{t}\}$, we write $S_{t}:=\sum_{s=1}^{t}\eta_{s}$ and $V_{t}:=\sum_{s=1}^{t}\eta_{s}^{2}$, and use $\bar{x}_{t}:=S_{t}^{-1}\sum_{s=1}^{t}\eta_{s}x_{s}$ to denote the weighted average iterate. Throughout, $f^{\star}:=f(x^{\star})$ denotes the optimal value and $\bar{F}_{t}$ the weighted average suboptimality defined in (4). Probabilities and expectations are taken with respect to the randomness in the stochastic gradients.
2 Anytime-Valid Confidence Sequences
In this section we derive an adaptive, anytime-valid confidence sequence for the weighted average suboptimality of projected SGD. Our goal is to construct, for each confidence level $\alpha\in(0,1)$, an upper confidence sequence $\{U_{t}(\alpha)\}$ in the sense of Definition 1.3 such that $\{\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\}$ holds with probability $1-\alpha$, where $\bar{F}_{t}$ is the weighted suboptimality defined in (4). A central feature of our construction is adaptivity: the bound adjusts to the realized optimization trajectory through the distances $\|x_{t}-x^{\star}\|$ and the squared stochastic gradient norms $\|g_{t}\|^{2}$, rather than relying solely on global parameters such as $R_{x}$ and $G$. A simple corollary will later convert this adaptive bound into a fully observable version involving only $R_{x}$, the realized quantities $\|g_{t}\|^{2}$, and the stepsizes.
To state the main result, we introduce the squared distance process
$Z_{t}:=\|x_{t}-x^{\star}\|^{2},$
and a trajectory-dependent quantity that governs the width of the confidence sequence: for each $t\geq 1$, define
$\sigma_{t}^{2}:=\sigma^{2}\eta_{t}^{2}Z_{t},\qquad\Sigma_{t}^{2}:=\sum_{s=1}^{t}\sigma_{s}^{2},$ (8)
where $\sigma^{2}>0$ is the sub-Gaussian variance proxy appearing in Assumption 1.1(iii). The cumulative term $\Sigma_{t}^{2}$ will emerge as a key scale parameter in our time-uniform concentration bound: larger values correspond to trajectories with larger stepsizes or larger distances to the optimum, and therefore lead to wider confidence sequences. Although $Z_{t}$ is not observable, it is always bounded by $R_{x}^{2}$, and keeping it explicit highlights which components of the trajectory influence the final bound. A fully observable, worst-case version can be obtained by replacing $Z_{t}$ with $R_{x}^{2}$.
Our confidence sequence will be stated in terms of the following data-dependent upper boundary:
$U_{t}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\,\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (9)
where $C,c>0$ are constants, provided explicitly below. This boundary combines three sources of variability: the cumulative variance proxy $\Sigma_{t}^{2}$, the initial distance $Z_{1}$, and the accumulated squared gradient steps $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$. The leading term involving $\Sigma_{t}^{2}$ captures the stochastic fluctuations of the process and stems from a time-uniform concentration inequality applied to an adapted sequence satisfying conditional sub-Gaussian tail bounds (see the proof of Theorem 2.2). In contrast, the terms $Z_{1}$ and $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ come directly from the deterministic structure of projected SGD, and appear when the distance recursion is telescoped.
Before proceeding, we introduce a mild technical condition on $\Sigma_{t}^{2}$ that simplifies our analysis.
Assumption 2.1 (Minimal cumulative variance proxy).
Fix $\alpha\in(0,0.6)$. We assume
$\Sigma_{t}^{2}\geq 2\Big(\log\frac{\pi^{2}}{6\alpha}+1\Big)\qquad\forall t\geq 1.$
This assumption is practically unrestrictive. Since $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$, the quantity $\Sigma_{t}^{2}$ typically exceeds this threshold after only a few iterations; if not, one may simply slightly enlarge the first stepsize $\eta_{1}$. Moreover, we specify the constants appearing in $U_{t}(\alpha)$:
$C\;:=\;6\sqrt{\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\}},\qquad c\coloneqq\log\frac{\pi^{2}}{6}.$ (10)
Note that under Assumption 1.2, $\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty}<\infty$, with $V_{\infty}\coloneqq\sum_{s=1}^{\infty}\eta_{s}^{2}$, ensuring that $C$ is finite. With Assumption 2.1 and the definition of $C$, the confidence sequence $U_{t}(\alpha)$ in (9) is now fully specified. The next theorem establishes that $U_{t}(\alpha)$ forms an anytime-valid upper confidence sequence for the weighted suboptimality process $\{\bar{F}_{t}\}_{t\geq 1}$.
Theorem 2.2 (Adaptive anytime-valid suboptimality bound).
Suppose Assumptions 1.1, 1.2, and 2.1 hold, and let $\sigma_{t}^{2}$ and $\Sigma_{t}^{2}$ be defined as in (8). Then, for every $\alpha\in(0,0.6)$, the process $\{U_{t}(\alpha)\}_{t\geq 1}$ defined in (9), with constant $C$ defined in (10), satisfies
$\mathbb{P}\!\left(\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\right)\;\geq\;1-\alpha,$
where $\bar{F}_{t}$ is the weighted suboptimality in (4). In particular, $\{U_{t}(\alpha)\}_{t\geq 1}$ is an anytime-valid upper confidence sequence for $\{\bar{F}_{t}\}_{t\geq 1}$ in the sense of Definition 1.3.
A noteworthy feature of the bound in Theorem 2.2 is its asymptotic decay. Under the stochastic approximation stepsize conditions $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$, we have $S_{t}\to\infty$ and $\Sigma_{t}^{2}$ uniformly bounded, since the trajectory remains in a compact set and $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$ is summable. Moreover, by Lemma A.1, the series $\sum_{s=1}^{\infty}\eta_{s}^{2}\|g_{s}\|^{2}$ converges almost surely to a finite random limit. Consequently, we have the following result.
Proposition 2.3 (Asymptotic decay of the confidence sequence).
Under the conditions of Theorem 2.2, for any fixed $\alpha\in(0,0.6)$,
$U_{t}(\alpha)\to 0\quad\text{almost surely as }t\to\infty.$
In particular, the confidence sequence is not only valid for all $t\geq 1$ but also shrinks to zero at the classical $1/S_{t}$ rate of stochastic approximation (up to a finite random constant).
This proposition ensures that projected SGD, equipped with our anytime-valid bound, provably converges to the optimum with a data-dependent, high-probability certificate of progress.
Remark 2.4 (Optimality of the $1/S_{t}$ rate).
The $1/S_{t}$ rate in Theorem 2.2 cannot, in general, be improved, even in a very simple one-dimensional deterministic setting. Fix any stepsizes $\{\eta_{t}\}_{t\geq 1}$ with $\eta_{t}>0$, $\sum_{t}\eta_{t}=\infty$, and $\sum_{t}\eta_{t}^{2}<\infty$, and consider the quadratic model
$f(x):=\frac{\mu}{2}x^{2},\qquad\mu\in(0,1),$
with minimizer $x^{\star}=0$, $f^{\star}=0$, and deterministic oracle $g_{t}=\nabla f(x_{t})=\mu x_{t}$. Starting from $x_{1}=x_{0}\neq 0$, the projected SGD recursion reduces to the deterministic update $x_{t+1}=x_{t}-\eta_{t}g_{t}$, and the weighted cumulative suboptimality
$A_{t}:=\sum_{s=1}^{t}\eta_{s}\Big(f(x_{s})-f^{\star}\Big)$
is nondecreasing with
[…]
For any anytime-valid upper confidence sequence $\{U_{t}(\alpha)\}_{t\geq 1}$ over a model class containing this example, validity on this deterministic path forces $A_{t}\leq S_{t}U_{t}(\alpha)$ for all $t$, and hence
$S_{t}U_{t}(\alpha)\;\geq\;A_{1}\qquad\text{for all }t\geq 1.$
Thus, even in this benign one-dimensional model, no anytime-valid confidence sequence can shrink uniformly faster than order $1/S_{t}$. This example illustrates that the $1/S_{t}$ rate in our upper bound is essentially optimal from a time-uniform perspective. $\clubsuit$
Before turning to the proof, we record an immediate consequence of Theorem 2.2. The theorem itself produces an adaptive confidence sequence that depends on the (unobservable) distances $\|x_{t}-x^{\star}\|$ and on the realized stochastic gradients. In practice, one often wishes to work with fully observable bounds, expressed only in terms of the known constraint radius $R_{x}$, the realized squared gradient norms $\|g_{t}\|^{2}$, and the predictable stepsizes $\{\eta_{t}\}$. The following corollary provides such a variant.
Corollary 2.5 (Observable anytime-valid suboptimality bound).
Under the conditions of Theorem 2.2, suppose in addition that $\mathcal{X}$ is known to satisfy $\|x-x^{\star}\|\leq R_{x}$ for all $x\in\mathcal{X}$. Define
$\widetilde{\sigma}_{t}^{2}:=\sigma^{2}\eta_{t}^{2}R_{x}^{2},\qquad\widetilde{\Sigma}_{t}^{2}:=\sum_{s=1}^{t}\widetilde{\sigma}_{s}^{2}=\sigma^{2}R_{x}^{2}V_{t}.$
Then for every $\alpha\in(0,0.6)$ the process
$U_{t}^{\mathrm{obs}}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\sqrt{\widetilde{\Sigma}_{t}^{2}\!\left(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\widetilde{\Sigma}_{t}^{2})\right)}+R_{x}^{2}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg)$
is an anytime-valid upper confidence sequence for $\{\bar{F}_{t}\}_{t\geq 1}$.
We highlight that for Corollary 2.5, the technical Assumption 2.1 need only be imposed on the observable variance proxy $\widetilde{\Sigma}_{t}^{2}$. Note that $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable and can be computed online at each iteration using only observable information.
2.1 Proof of Theorem 2.2
In this subsection, we unpack the proof of Theorem 2.2 through a sequence of conceptual steps. Rather than presenting all technical details at once, we isolate the key ideas that transform the projected SGD recursion into an anytime-valid confidence sequence. We adopt this layered presentation, along with detailed explanations, deliberately for its didactic value. We first construct a process whose one-step evolution compares the actual suboptimality to what is explained by the geometry of projected SGD and the stochastic gradients, and show that this process enjoys a trajectory-dependent sub-Gaussian control. We then present a time-uniform concentration lemma for martingale sums, after which a simple telescoping argument and normalization by $S_{t}$ yield the desired bound on $\bar{F}_{t}$.
Step 1: Constructing a sub-Gaussian process
We begin by identifying a natural potential function that tracks the progress of projected SGD toward the minimizer. A standard choice in convex optimization is the squared distance to the optimum, $Z_{t}=\|x_{t}-x^{\star}\|^{2}$. Because the update $x_{t+1}=\Pi_{\mathcal{X}}(x_{t}-\eta_{t}g_{t})$ is a Euclidean projection onto a closed convex set, the evolution of $Z_{t}$ can be controlled using the non-expansiveness of projection:
$Z_{t+1}\;=\;\|x_{t+1}-x^{\star}\|^{2}\;\leq\;\|x_{t}-\eta_{t}g_{t}-x^{\star}\|^{2}\;=\;Z_{t}-2\eta_{t}\langle x_{t}-x^{\star},g_{t}\rangle+\eta_{t}^{2}\|g_{t}\|^{2}.$ (11)
Rearranging this gives the pathwise lower bound $Z_{t}-Z_{t+1}\geq 2\eta_{t}\langle x_{t}-x^{\star},g_{t}\rangle-\eta_{t}^{2}\|g_{t}\|^{2}$. To connect this distance recursion to the suboptimality we care about, we package the suboptimality, the distance decrement, and the squared gradient step into a single quantity
$X_{t}\;:=\;2\eta_{t}\big(f(x_{t})-f(x^{\star})\big)-(Z_{t}-Z_{t+1})-\eta_{t}^{2}\|g_{t}\|^{2},$ (12)
with $X_{0}\coloneqq 0$. The next lemma shows that $X_{t}$ has two crucial properties: it has nonpositive conditional drift, and its conditional moment generating function is dominated by that of a sub-Gaussian variable with variance proxy $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$, as defined in (8). This is where convexity of $f$, the unbiasedness assumption, and the conditional sub-Gaussian noise assumption enter.
Lemma 2.6 (Conditional sub-Gaussian control).
Let $X_{t}$ be defined as in (12). Under Assumption 1.1, the following hold for all $t\geq 1$:
  • (i)
    Drift negativity:
    $\mathbb{E}[X_{t}\mid\mathcal{F}_{t-1}]\;\leq\;0.$ (13)
  • (ii)
    Conditional sub-Gaussian bound: for all $\lambda\geq 0$,
    $\mathbb{E}\!\left[\exp\big(\lambda X_{t}\big)\,\Big|\,\mathcal{F}_{t-1}\right]\;\leq\;\exp\!\big(2\lambda^{2}\sigma_{t}^{2}\big).$ (14)
Proof.
We first derive a convenient upper bound on $X_{t}$. From (11) and the definition (12),
$X_{t}$ $=2\eta_{t}\big(f(x_{t})-f(x^{\star})\big)-(Z_{t}-Z_{t+1})-\eta_{t}^{2}\|g_{t}\|^{2}$
$\leq 2\eta_{t}\big(f(x_{t})-f(x^{\star})\big)-\big(2\eta_{t}\langle x_{t}-x^{\star},g_{t}\rangle-\eta_{t}^{2}\|g_{t}\|^{2}\big)-\eta_{t}^{2}\|g_{t}\|^{2}$
$=2\eta_{t}\Big(f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},g_{t}\rangle\Big).$ (15)
Recalling that $\xi_{t}:=g_{t}-\nabla f(x_{t})$, we can further decompose $\langle x_{t}-x^{\star},g_{t}\rangle=\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle+\langle x_{t}-x^{\star},\xi_{t}\rangle$. Substituting this into (15) yields
$X_{t}\leq 2\eta_{t}\Big(f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle\Big)-2\eta_{t}\langle x_{t}-x^{\star},\xi_{t}\rangle.$ (16)
We now prove assertion (i). By convexity of $f$,
$f(x_{t})-f(x^{\star})\leq\langle\nabla f(x_{t}),x_{t}-x^{\star}\rangle,$
so the first term in (16) is nonpositive: $f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle\leq 0$. Taking conditional expectations in (16) and using the unbiasedness in Assumption 1.1(i), we obtain
$\mathbb{E}[X_{t}\mid\mathcal{F}_{t-1}]$ $\leq 2\eta_{t}\,\mathbb{E}\!\left[f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle\,\Big|\,\mathcal{F}_{t-1}\right]-2\eta_{t}\,\mathbb{E}\!\left[\langle x_{t}-x^{\star},\xi_{t}\rangle\,\Big|\,\mathcal{F}_{t-1}\right]$
$=2\eta_{t}\Big(f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle\Big)-2\eta_{t}\langle x_{t}-x^{\star},\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t-1}]\rangle$
$=2\eta_{t}\Big(f(x_{t})-f(x^{\star})-\langle x_{t}-x^{\star},\nabla f(x_{t})\rangle\Big)\;\leq\;0,$
which proves (13).
We now prove assertion (ii). From (16) and the convexity inequality just used, we also have the sharper pathwise bound
$X_{t}\;\leq\;-2\eta_{t}\langle x_{t}-x^{\star},\xi_{t}\rangle.$ (17)
Fix $\lambda\geq 0$. Multiplying (17) by $\lambda$ and taking conditional expectations,
$\mathbb{E}\!\left[\exp\big(\lambda X_{t}\big)\,\Big|\,\mathcal{F}_{t-1}\right]$ $\leq\mathbb{E}\!\left[\exp\!\big(-2\lambda\eta_{t}\langle x_{t}-x^{\star},\xi_{t}\rangle\big)\,\Big|\,\mathcal{F}_{t-1}\right].$
Applying Assumption 1.1(iii) with
$u:=-2\eta_{t}(x_{t}-x^{\star})\ \implies\ \|u\|^{2}=4\eta_{t}^{2}\|x_{t}-x^{\star}\|^{2}=4\eta_{t}^{2}Z_{t},$
we obtain
$\mathbb{E}\!\left[\exp\big(\lambda X_{t}\big)\,\Big|\,\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\sigma^{2}\|u\|^{2}}{2}\right)=\exp\!\big(2\lambda^{2}\sigma^{2}\eta_{t}^{2}Z_{t}\big)=\exp\!\big(2\lambda^{2}\sigma_{t}^{2}\big),$
which is exactly (14). ∎
The relevance of the process $X_{t}$ emerges once we consider its partial sums
$\bar{X}_{t}\coloneqq\sum_{s=1}^{t}X_{s},$ (18)
which admit a direct algebraic link to the cumulative weighted suboptimality:
$\bar{X}_{t}=2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)-(Z_{1}-Z_{t+1})-\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
Since $Z_{t+1}\geq 0$, it follows that
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\bar{X}_{t}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
Dividing by $2S_{t}$ and recalling the definition of $\bar{F}_{t}$ in (4), we obtain
$\bar{F}_{t}\;\leq\;\frac{1}{2S_{t}}\left(\bar{X}_{t}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\right).$
This inequality already mirrors the structure of the confidence sequence $U_{t}(\alpha)$ defined in (9): the first term comes from an upper bound on $\bar{X}_{t}$, while the remaining two terms carry over unchanged. Thus, controlling $\bar{X}_{t}$ uniformly over time is the key step in constructing $U_{t}(\alpha)$.
Step 2: A time-uniform concentration bound for $\bar{X}_{t}$
Our goal is now to show that the partial sums $\bar{X}_{t}$ admit a time-uniform upper bound in terms of the cumulative variance proxy $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma_{s}^{2}$. This is where the main probabilistic machinery enters: we construct an exponential supermartingale for each fixed $\lambda$, mix over a geometric grid of $\lambda$’s, and apply Ville’s inequality to obtain a high-probability bound that holds simultaneously for all $t$.
To state the result, define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\},$ (19)
with $\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty}<\infty$, and set $\lambda_{0}:=\sqrt{c_{1}/\sqrt{e}}$. Moreover, for $k=0,1,2,\ldots$, define
$\lambda_{k}:=\lambda_{0}e^{-k/2},\qquad w_{k}:=\frac{6}{\pi^{2}(k+1)^{2}}.$
Note that $\sum_{k\geq 0}w_{k}=1$. We are now ready to state the time-uniform concentration result for $\bar{X}_{t}$.
Lemma 2.7 (Time-uniform bound for $\bar{X}_{t}$).
Fix $\alpha\in(0,0.6)$, and let $\{X_{t}\}_{t\geq 0}$ be the process defined in (12). For $t\geq 0$, define the mixture process
$\bar{E}_{t}:=\sum_{k\geq 0}w_{k}\exp\!\left(\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\,\Sigma_{t}^{2}\right),$
with $\Sigma_{t}^{2}$ defined in (8). Then $\{\bar{E}_{t}\}_{t\geq 0}$ is a nonnegative supermartingale with $\bar{E}_{0}=1$. Moreover, on the event $\mathcal{E}_{\alpha}:=\bigl\{\forall t\geq 1:\bar{E}_{t}<\tfrac{1}{\alpha}\bigr\}$, which satisfies $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$, the following holds:
$\forall t\geq 1:\ \bar{X}_{t}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\frac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}.$
with constants $C,c$ defined in (10).
Proof.
For clarity, we divide the proof into six steps.
Step (i): $\{\bar{E}_{t}\}_{t\geq 0}$ is a nonnegative supermartingale. For any fixed $\lambda\geq 0$, define
$E_{t}(\lambda):=\exp\!\left(\lambda\bar{X}_{t}-2\lambda^{2}\Sigma_{t}^{2}\right)\qquad\forall t\geq 0,$
with the convention $\bar{X}_{0}=0$ and $\Sigma_{0}^{2}=0$, so $E_{0}(\lambda)=1$. Using $\bar{X}_{t}=\bar{X}_{t-1}+X_{t}$ and $\Sigma_{t}^{2}=\Sigma_{t-1}^{2}+\sigma_{t}^{2}$, we compute
$\mathbb{E}[E_{t}(\lambda)\mid\mathcal{F}_{t-1}]$ $=\mathbb{E}\!\left[\exp\!\left(\lambda(\bar{X}_{t-1}+X_{t})-2\lambda^{2}(\Sigma_{t-1}^{2}+\sigma_{t}^{2})\right)\Bigm|\mathcal{F}_{t-1}\right]$
$=\exp\!\left(\lambda\bar{X}_{t-1}-2\lambda^{2}\Sigma_{t-1}^{2}\right)\mathbb{E}\!\left[\exp\!\left(\lambda X_{t}-2\lambda^{2}\sigma_{t}^{2}\right)\Bigm|\mathcal{F}_{t-1}\right]$
$=E_{t-1}(\lambda)\,\mathbb{E}\!\left[\exp\!\left(\lambda X_{t}-2\lambda^{2}\sigma_{t}^{2}\right)\Bigm|\mathcal{F}_{t-1}\right].$
By the conditional sub-Gaussian assumption (14), the conditional expectation in the last line is at most $1$. Hence
$\mathbb{E}[E_{t}(\lambda)\mid\mathcal{F}_{t-1}]\leq E_{t-1}(\lambda),$
so $\{E_{t}(\lambda)\}_{t\geq 0}$ is a nonnegative supermartingale with $E_{0}(\lambda)=1$. Finally, since each $E_{t}(\lambda_{k})$ is a nonnegative supermartingale and $(w_{k})_{k\geq 0}$ is a fixed sequence of nonnegative weights summing to one, $\{\bar{E}_{t}\}_{t\geq 0}$ is also a nonnegative supermartingale with $\bar{E}_{0}=\sum_{k\geq 0}w_{k}=1$.
Step (ii): A first upper bound on $\bar{X}_{t}$. By Ville’s inequality applied to the nonnegative supermartingale $\bar{E}_{t}$,
$\mathbb{P}\!\left(\sup_{t\geq 1}\bar{E}_{t}\geq\tfrac{1}{\alpha}\right)\leq\alpha.$
Equivalently, the event
$\mathcal{E}_{\alpha}:=\Big\{\forall t\geq 1:\ \bar{E}_{t}<\tfrac{1}{\alpha}\Big\}$
satisfies $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$. Fix $\omega\in\mathcal{E}_{\alpha}$ and a time $t\geq 1$ with $\Sigma_{t}^{2}>0$. For notational simplicity, we suppress the dependence on $\omega$ in what follows. On $\mathcal{E}_{\alpha}$ we have
$\bar{E}_{t}=\sum_{k\geq 0}w_{k}\exp\!\left(\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\Sigma_{t}^{2}\right)<\frac{1}{\alpha}.$
Hence for every $k\geq 0$,
$w_{k}\exp\!\left(\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\Sigma_{t}^{2}\right)<\frac{1}{\alpha},$
which implies
$\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\Sigma_{t}^{2}<\log\frac{1}{\alpha w_{k}}.$
Rearranging this inequality yields, for each $k\geq 0$,
$\bar{X}_{t}<\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}\Sigma_{t}^{2}=:B_{k}(\Sigma_{t}^{2}).$
Since this holds simultaneously for all $k\geq 0$, we get the deterministic bound
$\bar{X}_{t}\leq\inf_{k\geq 0}B_{k}(\Sigma_{t}^{2}).$ (20)
It remains to upper bound the infimum in terms of $\Sigma_{t}^{2}$ and $\alpha$.
Step (iii): Explicit form of $B_{k}(\Sigma_{t}^{2})$. For notational simplicity, set
$v:=\Sigma_{t}^{2}>0,\qquad L_{\alpha}:=\log\frac{\pi^{2}}{6\alpha}.$
Using $w_{k}={6}/(\pi^{2}(k+1)^{2})$ and $\lambda_{k}=\lambda_{0}e^{-k/2}$, we have
$\log\frac{1}{\alpha w_{k}}=\log\frac{\pi^{2}}{6\alpha}+2\log(k+1)=L_{\alpha}+2\log(k+1),$
and
$\frac{1}{\lambda_{k}}=\frac{e^{k/2}}{\lambda_{0}}.$
Hence, for $v>0$,
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v=\frac{e^{k/2}}{\lambda_{0}}\big(L_{\alpha}+2\log(k+1)\big)+2\lambda_{0}e^{-k/2}v.$
Step (iv): Choose $k$ as a function of $v$. Define
$L_{v}:=\log\log(\mathrm{e}+v)>0.$
We will pick an integer $k=k(v)$ such that:
  1. 1.
    $e^{k(v)/2}$ is within a constant factor of
    $z_{v}:=\sqrt{\frac{v}{L_{\alpha}+L_{v}}},$
  2. 2.
    $\log(k(v)+1)$ is bounded by a constant multiple of $L_{v}$.
One convenient choice is
$k(v):=\left\lfloor\log\frac{v}{L_{\alpha}+L_{v}}\right\rfloor\qquad\text{for }v>0,$
which immediately implies that $e^{k(v)/2}\leq z_{v}\leq e^{(k(v)+1)/2}$. Assumption 2.1 and Lemma A.3 (note that $L_{\alpha}\geq 1$ holds since $\alpha\leq 0.6$) guarantee that $k(v)\geq 0$. Therefore, for this choice, we have that
$\frac{1}{\sqrt{e}}\,z_{v}\leq e^{k(v)/2}\leq z_{v},\qquad\log(k(v)+1)\leq c_{1}L_{v},$ (21)
with the constant $c_{1}\geq 1$ defined in (19). Note that the inequality $\log(k(v)+1)\leq c_{1}L_{v}$ follows from Lemma A.2, using the fact that $L_{\alpha}+L_{v}>L_{\alpha}$. For this $k=k(v)$, let
$D_{v}:=L_{\alpha}+2\log(k(v)+1).$
Using (21), we have $2\log(k(v)+1)\leq 2c_{1}L_{v}$, so
$D_{v}\leq L_{\alpha}+2c_{1}L_{v}\leq c_{2}\big(L_{\alpha}+L_{v}\big)$
for the constant $c_{2}\coloneqq\max\{1,2c_{1}\}=2c_{1}$.
Step (v): Bounding $B_{k(v)}(v)$. Using (21) again, we have
$e^{k(v)/2}\leq z_{v}=\sqrt{\frac{v}{L_{\alpha}+L_{v}}},\qquad e^{-k(v)/2}\leq\sqrt{e}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}.$
Therefore,
$\frac{e^{k(v)/2}}{\lambda_{0}}D_{v}\leq\frac{1}{\lambda_{0}}\sqrt{\frac{v}{L_{\alpha}+L_{v}}}\,c_{2}\big(L_{\alpha}+L_{v}\big)=\frac{c_{2}}{\lambda_{0}}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
and
$2\lambda_{0}e^{-k(v)/2}v\leq 2\lambda_{0}\sqrt{e}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}\,v=2\sqrt{e}\lambda_{0}\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
Summing these two bounds and recalling that $\lambda_{0}=\sqrt{{c_{2}}/({2\sqrt{e}})}$, we obtain
$B_{k(v)}(v)\leq 2\sqrt{2\sqrt{e}c_{2}}\sqrt{v\big(L_{\alpha}+L_{v}\big)}\leq c_{3}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
for the constant $c_{3}\coloneqq 4\sqrt{c_{2}}\leq 6\sqrt{c_{1}}=:C$.
Finally, since $\inf_{k}B_{k}(v)\leq B_{k(v)}(v)$ for each fixed $v>0$, we conclude that for all $v>0$,
$\inf_{k\geq 0}B_{k}(v)\leq C\sqrt{v\big(L_{\alpha}+\log\log(\mathrm{e}+v)\big)}=C\sqrt{v\big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+v)\big)},$
where $c\coloneqq\log\tfrac{\pi^{2}}{6}$.
Step (vi): Conclude the time-uniform bound. Combining this with (20), we obtain, on the event $\mathcal{E}_{\alpha}$,
$\bar{X}_{t}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}\qquad\forall t\geq 1,$
which gives the stated bound. ∎
This lemma is the probabilistic engine of the paper: whenever one can certify a conditional sub-Gaussian inequality of the form (14), one obtains a time-uniform concentration inequality for the partial sums, with a variance scale $\Sigma_{t}^{2}$ and only a mild iterated-logarithm correction. In particular, $\bar{X}_{t}$ remains controlled simultaneously over all $t\geq 1$ with probability at least $1-\alpha$.
Step 3: From martingale sums to cumulative suboptimality.
Using the definition of $X_{t}$, we now rewrite $\bar{X}_{t}$ in terms of the cumulative weighted suboptimality, the distance terms $Z_{t}$, and the squared gradient terms. As explained at the end of Step 1, a simple telescoping argument yields the desired upper bound on the numerator of $\bar{F}_{t}$.
Lemma 2.8 (From $\bar{X}_{t}$ to cumulative suboptimality).
Let $X_{t}$ be defined as in (12), and let $\sigma_{t}^{2}$ and $\Sigma_{t}^{2}$ be as in (8). Then, for any $\alpha\in(0,0.6)$, with probability at least $1-\alpha$, the following holds simultaneously for all $t\geq 1$:
$\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\frac{1}{2}\Bigg(C\,\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (22)
where $C>0$ is the constant from Lemma 2.7.
Proof.
From Lemma 2.7 we know that for any $\alpha\in(0,0.6)$, there exists an event $\mathcal{E}_{\alpha}$ with $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$ such that, on $\mathcal{E}_{\alpha}$,
$\forall t\geq 1:\qquad\sum_{s=1}^{t}X_{s}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}.$ (23)
Next, we express $\sum_{s=1}^{t}X_{s}$ in terms of the cumulative suboptimality. Summing the definition (12) from $s=1$ to $t$ gives
$\sum_{s=1}^{t}X_{s}$ $=2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)-\sum_{s=1}^{t}(Z_{s}-Z_{s+1})-\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$
$=2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)-(Z_{1}-Z_{t+1})-\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
Rearranging this identity, we obtain
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)=\sum_{s=1}^{t}X_{s}+(Z_{1}-Z_{t+1})+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$ (24)
On the event $\mathcal{E}_{\alpha}$, we may bound $\sum_{s=1}^{t}X_{s}$ using (23). Moreover, $Z_{t+1}=\|x_{t+1}-x^{\star}\|^{2}\geq 0$, so we can drop it. Plugging these bounds into (24), we find that on $\mathcal{E}_{\alpha}$, for all $t\geq 1$,
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
Dividing both sides by $2$ yields (22), and the bound holds simultaneously for all $t\geq 1$ on the event $\mathcal{E}_{\alpha}$, which has probability at least $1-\alpha$. ∎
This lemma is the bridge from the martingale world back to optimization: it translates the time-uniform control of the process $\bar{X}_{t}$ into a time-uniform upper bound on the cumulative weighted suboptimality $\sum_{s=1}^{t}\eta_{s}(f(x_{s})-f^{\star})$. Finally, dividing both sides by $S_{t}$ yields a time-uniform bound for the weighted average suboptimality $\bar{F}_{t}$, completing the proof of Theorem 2.2.
2.2 Clarifying remarks and intuition
The proof of Theorem 2.2 relies on several ideas from time-uniform sequential analysis that may be unfamiliar to readers with a background in stochastic optimization. In this subsection we collect three clarifying remarks intended to illuminate the construction. First, we explain the role of the mixture supermartingale $\bar{E}_{t}$ and why a single-parameter exponential martingale cannot yield a time-uniform bound without sacrificing optional stopping. Second, we provide a heuristic “continuous optimization” argument that motivates the geometric grid over $\lambda_{k}$ and the emergence of an iterated-logarithm term in the boundary. Third, we discuss why our bound does not contain a $\log\log t$ factor typical in classical anytime concentration results, and how the square-summable stepsize condition fundamentally alters the variance accumulation in the stochastic approximation setting.
Remark 2.9 (Why the mixture $\bar{E}_{t}$ is necessary).
To understand the need for mixtures in Lemma 2.7, it is instructive to redo the key step with a single exponential supermartingale. For any fixed $\lambda>0$, define
$E_{t}(\lambda):=\exp\!\big(\lambda\bar{X}_{t}-2\lambda^{2}\Sigma_{t}^{2}\big),$
as in Step (i) of Lemma 2.7. By the conditional sub-Gaussian property (14), $\{E_{t}(\lambda)\}_{t\geq 0}$ is a nonnegative supermartingale with $E_{0}(\lambda)=1$. Ville’s inequality therefore gives
$\mathbb{P}\!\left(\sup_{t\geq 1}E_{t}(\lambda)\,\geq\,\frac{1}{\alpha}\right)\,\leq\,\alpha.$ (25)
Equivalently, with probability at least $1-\alpha$,
$\forall t\geq 1:\ E_{t}(\lambda)<\frac{1}{\alpha}\ \iff\ \lambda\bar{X}_{t}-2\lambda^{2}\Sigma_{t}^{2}<\log\frac{1}{\alpha}.$
Rearranging yields the deterministic upper bound
$\bar{X}_{t}<\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t}^{2}\qquad\forall t\geq 1.$ (26)
Now fix a time $t$ and, for that $t$, treat the right-hand side as a function of $\lambda$:
$B(\lambda,t):=\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t}^{2},\qquad\lambda>0.$ (27)
This function is minimized explicitly at
$\lambda^{\star}(t)=\sqrt{\frac{\log(1/\alpha)}{2\Sigma_{t}^{2}}},$ (28)
with minimum value
$B\big(\lambda^{\star}(t),t\big)=2\sqrt{2}\sqrt{\Sigma_{t}^{2}\log\frac{1}{\alpha}}.$ (29)
Thus, for a fixed time $t$, a single exponential supermartingale produces an upper bound with the scaling $\sqrt{\Sigma_{t}^{2}\log(1/\alpha)}$. This matches the leading term that appears in Theorem 2.2, but without the additional iterated-logarithm factor that ultimately arises from restricting $\lambda$ to lie on a fixed grid.
The problem arises from the fact that $\lambda^{\star}(t)$ in (28) depends on $\Sigma_{t}^{2}$ and is therefore $\mathcal{F}_{t}$-measurable: it is chosen after observing the data up to time $t$. Substituting this data-dependent $\lambda^{\star}(t)$ back into $E_{t}(\lambda)$ destroys the supermartingale structure, so Ville’s inequality cannot be applied time-uniformly. With a single exponential process one faces a choice: either fix $\lambda$ in advance (yielding time-uniform but suboptimal bounds for many $t$), or optimize $\lambda$ at each $t$ (yielding sharp fixed-time bounds but no optional-stopping guarantee).
The mixture construction in Lemma 2.7 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [de2009self]). Modern confidence-sequence constructions, such as the sub-$\psi$ mixture framework in [howard2020time, howard2021time] and the e-process approach in [waudby2024estimating], rely crucially on this principle. $\clubsuit$
Remark 2.10 (Intuition for the geometric grid $\{\lambda_{k}\}_{k\geq 0}$).
The mixture construction in Lemma 2.7 hinges on choosing a geometric grid of tuning parameters $\lambda_{k}=\lambda_{0}e^{-k/2}$. Here we provide intuition for this choice by examining how the envelope $B_{k}(\Sigma_{t}^{2})$ is minimized, and how the geometry of the grid matches the natural scale of the problem. Fix $v>0$ and recall the upper envelope
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v,\qquad\lambda_{k}=\lambda_{0}e^{-k/2},\quad w_{k}=\frac{6}{\pi^{2}(k+1)^{2}}.$
Momentarily regard $k$ as a continuous variable and write $z=e^{k/2}>0$. Then
$B_{k}(v)=\frac{D(z)}{\lambda_{0}}\,z+2\lambda_{0}\,\frac{v}{z},$
where
$D(z):=\log\frac{\pi^{2}}{6\alpha}+2\log(k+1)=\log\frac{\pi^{2}}{6\alpha}+2\log\!\Big(1+2\log z\Big).$
The key observation is that $D(z)$ varies very slowly in $z$: once $z$ is chosen near its optimal scale, $D(z)$ grows only like $\log\log(\mathrm{e}+z)$, a negligible rate compared to any power of $z$. As a heuristic, we therefore treat $D(z)$ as approximately constant and consider the simplified objective
$B(z,v):=\frac{D}{\lambda_{0}}\,z+2\lambda_{0}\,\frac{v}{z},$
where $D>0$ is held fixed. The function $B(z,v)$ is strictly convex in $z>0$, and its minimum is attained at
$z^{\star}=\sqrt{\frac{2\lambda_{0}^{2}v}{D}}\ \implies\ \lambda^{\star}=\frac{\lambda_{0}}{z^{\star}}=\sqrt{\frac{D}{2v}}.$
Substituting $z^{\star}$ back into $B(z,v)$ yields
$\min_{z>0}B(z,v)=2\sqrt{2Dv}.$
Heuristically, since $D(z)\approx L_{\alpha}+\log\log(\mathrm{e}+z)\approx L_{\alpha}+\log\log(\mathrm{e}+v)$ at the relevant scale, this suggests the leading-order behavior
$\inf_{k\geq 0}B_{k}(v)\;\lesssim\;\sqrt{v\Big(L_{\alpha}+\log\log(\mathrm{e}+v)\Big)}.$
The slowly varying term $\log(k+1)$ introduces an iterated-logarithm correction, where the $\log\log(\mathrm{e}+v)$ factor arises precisely from the slow growth of $\log(k+1)$ on this geometric grid. In particular, for a fixed time $t$, the optimal value of $\lambda$ is of order $v^{-1/2}$, where $v=\Sigma_{t}^{2}$ controls the variance scale of the martingale sum. Since $v$ can vary over many orders of magnitude as $t$ increases, a geometric grid of the form $\lambda_{k}=\lambda_{0}e^{-k/2}$ is a natural discretization: consecutive grid points differ by a constant factor, ensuring that at least one $\lambda_{k}$ lies within a constant factor of the ideal $\lambda^{\star}\approx v^{-1/2}$, for every $t$.
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This type of “scale-matching” intuition for geometric grids is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion preceding Theorem 1 in [howard2021time] and the grid-selection heuristics in the sub-$\psi$ mixture framework in [howard2020time]. $\clubsuit$
Remark 2.11 (Why no $\log\log t$ term appears).
A recurring feature in anytime martingale bounds is the presence of iterated logarithm terms such as $\log\log t$ or $\log\log d$ (see, e.g., Robbins–Siegmund iterated LIL bounds and mixture inequalities in [howard2020time]). At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t}^{2}\log\log t}$ rather than the $\sqrt{\Sigma_{t}^{2}\log\log(\mathrm{e}+\Sigma_{t}^{2})}$ term appearing in Theorem 2.2. The difference comes from the square-summable stepsize assumption in Assumption 1.2 and the resulting definition of $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$. In most sequential testing settings, $\Sigma_{t}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.7 ultimately leads to iterated-logarithm terms in $\log t$. Here, however, $\{\eta_{s}\}_{s\geq 1}$ is square summable, so $\Sigma_{t}^{2}$ is bounded and converges to
$\Sigma_{\infty}^{2}:=\sum_{s=1}^{\infty}\sigma^{2}\eta_{s}^{2}Z_{s}\;<\;\infty\qquad\text{almost surely},$
under Assumption 1.2. Consequently, the relevant “scale” for the mixture optimization is the value of $\Sigma_{t}^{2}$ itself, not the time index $t$. This is why the iterated-logarithm term takes the form $\log\log(\mathrm{e}+\Sigma_{t}^{2})$ rather than $\log\log t$: the penalty adapts to the variance process instead of time. Because $\sum_{t\geq 1}\eta_{t}^{2}<\infty$ ensures that $\Sigma_{t}^{2}$ is almost surely bounded, this “variance-based iterated log” is substantially smaller than the classical “time-based iterated log” that arises when $\Sigma_{t}^{2}$ grows linearly with $t$.
This phenomenon is well known in martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (predictable quadratic variation, conditional variance processes, or self-normalized sums) shrink faster than those written in terms of time alone; see [de2009self] and the “variance process versus time” discussion in [howard2021time, Section 3]. $\clubsuit$
3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn projected SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (i.e., they monitor validation loss, gradient norms, or empirical convergence patterns) and their guarantees, when available, typically hold only for a pre-specified time or for a specific deterministic stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies. Throughout this section, we work with the observable confidence sequence $U_{t}^{\mathrm{obs}}(\alpha)$, and Assumption 2.1 can be imposed on the observable variance proxy $\widetilde{\Sigma}_{t}^{2}=\sigma^{2}R_{x}^{2}V_{t}$.
Corollary 2.5 provides an anytime-valid upper confidence sequence $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ in (4). Combining this with the convexity bound $f(\bar{x}_{t})-f^{\star}\leq\bar{F}_{t}$, we obtain, for any fixed $\alpha\in(0,1)$,
$\mathbb{P}\!\left(\forall t\geq 1:\ f(\bar{x}_{t})-f^{\star}\leq U_{t}^{\mathrm{obs}}(\alpha)\right)\;\geq\;1-\alpha,$ (30)
where $\bar{x}_{t}=S_{t}^{-1}\sum_{s=1}^{t}\eta_{s}x_{s}$ is the weighted average iterate, with $S_{t}=\sum_{s=1}^{t}\eta_{s}$.
Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,1)$, it is therefore natural to stop as soon as the certificate $U_{t}^{\mathrm{obs}}(\alpha)$ falls below $\varepsilon$. As already foreshadowed in the introduction, we formalize this through the stopping time
$\tau_{\varepsilon}:=\inf\big\{t\geq 1:U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon\big\},$ (31)
with the convention $\inf\emptyset=\infty$. Since $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable for each $t$, $\tau_{\varepsilon}$ is a proper stopping time with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geq 1}$ generated by the iterates and gradients.
In the remainder of this section we show that, under the stochastic approximation conditions on the stepsize in Assumption 1.2, the stopping time $\tau_{\varepsilon}$ enjoys two key properties: (i) it is almost surely finite for any fixed $\varepsilon>0$, and (ii) the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ is $\varepsilon$-optimal with confidence $1-\alpha$. Together, these facts certify projected SGD as an anytime-valid $\varepsilon$-optimization algorithm driven entirely by observable quantities along the run.
Theorem 3.1 (Adaptive anytime-valid stopping rule).
Suppose Assumptions 1.1, 1.2, and 2.1 (with $\Sigma_{t}^{2}$ replaced by $\widetilde{\Sigma}_{t}^{2}$) hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.5. Fix $\varepsilon>0$ and $\alpha\in(0,0.6)$, and define the stopping time $\tau_{\varepsilon}$ by (31). Then:
  1. (i)
    $\tau_{\varepsilon}<\infty$ almost surely;
  2. (ii)
    the stopped iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies:
    $\mathbb{P}\Big(f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\Big)\;\geq\;1-\alpha.$ (32)
In particular, projected SGD equipped with $\tau_{\varepsilon}$ yields a data-dependent stopping strategy that guarantees $\varepsilon$-optimality with probability $1-\alpha$, without requiring a predetermined time horizon.
Proof.
We first show that $\tau_{\varepsilon}$ is almost surely finite. By Proposition 2.3, for any fixed $\alpha\in(0,0.6)$, we have
$U_{t}^{\mathrm{obs}}(\alpha)\to 0\quad\text{almost surely as }t\to\infty.$
Let
$\Omega_{0}:=\Big\{\lim_{t\to\infty}U_{t}^{\mathrm{obs}}(\alpha)=0\Big\},$
so that $\mathbb{P}(\Omega_{0})=1$. Fix $\varepsilon>0$ and $\omega\in\Omega_{0}$. By the definition of convergence, there exists a (random, finite) index $T_{\varepsilon}(\omega)\in\mathbb{N}$ such that
$U_{t}^{\mathrm{obs}}(\alpha)(\omega)\;\leq\;\varepsilon\qquad\forall t\geq T_{\varepsilon}(\omega).$
By the definition (31) of $\tau_{\varepsilon}$,
$\tau_{\varepsilon}(\omega)=\inf\big\{t\geq 1:[U_{t}^{\mathrm{obs}}(\alpha)](\omega)\leq\varepsilon\big\}\;\leq\;T_{\varepsilon}(\omega)\;<\;\infty.$
Since this holds for every $\omega\in\Omega_{0}$ and $\mathbb{P}(\Omega_{0})=1$, we conclude that $\tau_{\varepsilon}<\infty$ almost surely.
For part (ii), let
$\mathcal{E}_{\alpha}:=\Big\{\forall t\geq 1:\ f(\bar{x}_{t})-f^{\star}\leq U_{t}^{\mathrm{obs}}(\alpha)\Big\}.$
By (30), $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$. On the event $\mathcal{E}_{\alpha}$, the inequality $f(\bar{x}_{t})-f^{\star}\leq U_{t}^{\mathrm{obs}}(\alpha)$ holds for all $t\geq 1$. In particular, on $\mathcal{E}_{\alpha}$ we have
$f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\;\leq\;U_{\tau_{\varepsilon}}^{\mathrm{obs}}(\alpha).$
By definition of $\tau_{\varepsilon}$, $U_{\tau_{\varepsilon}}^{\mathrm{obs}}(\alpha)\leq\varepsilon$, so on $\mathcal{E}_{\alpha}$,
$f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\;\leq\;\varepsilon.$
Thus $\mathcal{E}_{\alpha}\subseteq\{f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\}$, and taking probabilities yields
$\mathbb{P}\Big(f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\Big)\;\geq\;\mathbb{P}(\mathcal{E}_{\alpha})\;\geq\;1-\alpha,$
which is exactly (32). This proves part (ii) and completes the proof. ∎
Theorem 3.1 shows that projected SGD equipped with the stopping rule $\tau_{\varepsilon}$ is a data-dependent $\varepsilon$-optimization procedure: one may monitor $U_{t}^{\mathrm{obs}}(\alpha)$ online and stop as soon as it falls below $\varepsilon$, without specifying a horizon in advance or invalidating the $1-\alpha$ guarantee. This yields fully implementable stopping rules based only on running sums of $\eta_{t}$, $\eta_{t}^{2}$, and the observed gradient norms, with the expected tradeoff between adaptivity and reliance on global parameters.
In the following subsection we study the stopping-time complexity associated with the rule $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$. We present explicit bounds with deterministic constants via a strengthening of our assumptions on the stochastic gradients.
3.1 Stopping-time complexity
In this subsection we derive explicit bounds on the stopping time $\tau_{\varepsilon}$ under standard stepsize schedules for stochastic approximation. Our assumptions on the stochastic gradients are as follows.
Assumption 3.2 (Stochastic gradients (strengthened)).
For all $t\geq 1$, the stochastic gradient $g_{t}$ satisfies:
  • (i)
    Unbiasedness: $\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}]=\nabla f(x_{t})$, and
  • (ii)
    Boundedness: $\|g_{t}\|\leq G$ almost surely.
Note that Assumption 3.2(ii) implies Assumption 1.1(ii), and moreover implies the conditional sub-Gaussian Assumption 1.1(iii) with variance proxy $\sigma^{2}:=4G^{2}$. Indeed, for any $\lambda\in\mathbb{R}$ and $u\in\mathbb{R}^{d}$,
$|\big\langle g_{t}-\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}],u\big\rangle|\leq 2G\|u\|\qquad\text{almost surely},$
so by Hoeffding’s lemma,
$\mathbb{E}\!\left[\exp\!\left(\lambda\,\langle g_{t}-\mathbb{E}[g_{t}\mid\mathcal{F}_{t-1}],u\rangle\right)\Bigm|\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\,4G^{2}\,\|u\|^{2}}{2}\right),$
establishing conditional sub-Gaussianity.
Before quantifying the dependence of $\tau_{\varepsilon}$ on the target accuracy $\varepsilon$, it is convenient to reduce the stopping criterion $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$ to a simple threshold condition on the accumulated stepsizes $S_{t}$. The following corollary makes this reduction precise.
Corollary 3.3 (General certified stopping bound).
Let Assumptions 3.2, 1.2, and 2.1 (with $\Sigma_{t}^{2}$ replaced by $\widetilde{\Sigma}_{t}^{2}$) hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.5. Moreover, define
$K_{\alpha}:=\frac{1}{2}\Bigg(C\sqrt{\widetilde{\Sigma}_{\infty}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log\!\big(\mathrm{e}+\widetilde{\Sigma}_{\infty}^{2}\big)\Big)}+R_{x}^{2}+G^{2}V_{\infty}\Bigg),$ (33)
with constant $C$ defined in (10), $\widetilde{\Sigma}_{\infty}^{2}:=4G^{2}R_{x}^{2}V_{\infty}$, and $V_{\infty}:=\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Then, for every target accuracy $\varepsilon>0$,
$\tau_{\varepsilon}\;\leq\;\inf\!\left\{t\geq 1:\;S_{t}\geq\frac{K_{\alpha}}{\varepsilon}\right\}\qquad\text{almost surely}.$
Corollary 3.3 shows that the stopping time $\tau_{\varepsilon}$ is controlled entirely by the growth of $S_{t}$: any deterministic lower bound on $S_{t}$ immediately yields an almost-sure upper bound on $\tau_{\varepsilon}$. The constant $K_{\alpha}$ depends only on the global problem parameters $(G,R_{x})$, the confidence level $\alpha$, and the stepsize schedule $\{\eta_{t}\}_{t\geq 1}$. In what follows, we apply Corollary 3.3 to several standard choices of stepsizes to obtain certified stopping-time complexity guarantees.
(1) Polynomially decaying stepsizes. Consider
$\eta_{t}=\eta_{0}t^{-\gamma},\qquad\eta_{0}>0,\ \ \gamma\in\Big(\frac{1}{2},1\Big).$ (34)
Since $\gamma>1/2$, this family satisfies $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Lemma A.4 implies that for all $t\geq 1$,
$S_{t}\geq\frac{\eta_{0}}{1-\gamma}\Big(t^{1-\gamma}-1\Big).$
Using Corollary 3.3, it suffices that
$\frac{\eta_{0}}{1-\gamma}\Big(t^{1-\gamma}-1\Big)\ \geq\ \frac{K_{\alpha}}{\varepsilon}.$
Solving for $t$ yields:
$\tau_{\varepsilon}\ \leq\ \Biggl(1+\frac{(1-\gamma)K_{\alpha}}{\eta_{0}\,\varepsilon}\Biggr)^{\!\!\frac{1}{1-\gamma}}\qquad\text{almost surely}.$ (35)
We summarize this in the following proposition.
Proposition 3.4 (Certified stopping complexity for $\gamma\in(1/2,1)$).
Under the conditions of Corollary 3.3 and for the stepsize schedule (34), the stopping time $\tau_{\varepsilon}$ defined in (31) satisfies
$\tau_{\varepsilon}=O\!\Big(\varepsilon^{-\frac{1}{1-\gamma}}\Big)\qquad\text{almost surely}.$
In particular, $\tau_{\varepsilon}$ scales polynomially in $1/\varepsilon$ with an exponent determined by the decay rate $\gamma$.
(2) Harmonic stepsizes. Now consider the boundary case
$\eta_{t}=\frac{\eta_{0}}{t},\qquad\eta_{0}>0.$ (36)
Then $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$, so Assumption 1.2 holds. Lemma A.5 gives $S_{t}\geq\eta_{0}\log t$. By Corollary 3.3, it suffices that
$\eta_{0}\log t\ \geq\ \frac{K_{\alpha}}{\varepsilon},$
which implies
$\tau_{\varepsilon}\ \leq\ \exp\!\Big(\frac{K_{\alpha}}{\eta_{0}\,\varepsilon}\Big)\qquad\text{almost surely}.$ (37)
We summarize this in the following proposition.
Proposition 3.5 (Certified stopping complexity for $\gamma=1$).
Under the conditions of Corollary 3.3 and for the stepsize schedule (36), the stopping time $\tau_{\varepsilon}$ defined in (31) satisfies
$\tau_{\varepsilon}=\exp\!\Big(O(\varepsilon^{-1})\Big)\qquad\text{almost surely}.$
Thus harmonic stepsizes yield certified stopping but with very slow complexity in $1/\varepsilon$.
Remark 3.6 (Practical implications).
The polynomial regime $\gamma\in(1/2,1)$ produces certified complexity $\tau_{\varepsilon}=O\Big(\varepsilon^{-1/(1-\gamma)}\Big)$, which matches the classical stochastic approximation tradeoff between stepsize decay and variance control. The harmonic schedule $\eta_{t}=\eta_{0}/t$, although valid for confidence-sequence stopping, leads to exponential complexity in $1/\varepsilon$ and is therefore unattractive in practice. Combined with Theorem 3.1, these results show that the stepsize schedule entirely governs the certified stopping rate, and that the confidence-sequence approach preserves standard stochastic approximation complexity guarantees while remaining valid under arbitrary, data-dependent optional stopping. $\clubsuit$
Acknowledgements
Liviu Aolaritei acknowledges support from the Swiss National Science Foundation through the Postdoc.Mobility Fellowship (grant agreement P500PT_222215). Funded in part by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Technical lemmas
Lemma A.1.
Let Assumptions 1.2 and  1.1(ii) hold. Then,
$\sum_{t=1}^{\infty}\eta_{t}^{2}\|g_{t}\|^{2}<\infty\qquad\text{almost surely}.$
Proof.
Define the nonnegative partial sums
$Y_{n}:=\sum_{t=1}^{n}\eta_{t}^{2}\|g_{t}\|^{2}\qquad\forall n\geq 1,$
and let $Y:=\sum_{t=1}^{\infty}\eta_{t}^{2}\|g_{t}\|^{2}:=\lim_{n\to\infty}Y_{n}$ (possibly taking values in $0,\infty]$). Since the sequence $\{Y_{n}\}_{n\geq 1}$ is nondecreasing and nonnegative, we may apply the monotone convergence theorem:
$\mathbb{E}[Y]=\mathbb{E}\Big[\lim_{n\to\infty}Y_{n}\Big]=\lim_{n\to\infty}\mathbb{E}[Y_{n}].$
For each $n$, we have
$\mathbb{E}[Y_{n}]=\sum_{t=1}^{n}\eta_{t}^{2}\,\mathbb{E}[\|g_{t}\|^{2}].$
Using the tower property and Assumption 1.1(ii),
$\mathbb{E}[\|g_{t}\|^{2}]=\mathbb{E}\big[\mathbb{E}[\|g_{t}\|^{2}\mid\mathcal{F}_{t-1}]\big]\leq\mathbb{E}[G^{2}]=G^{2},$
so that
$\mathbb{E}[Y_{n}]\leq\sum_{t=1}^{n}\eta_{t}^{2}G^{2}=G^{2}\sum_{t=1}^{n}\eta_{t}^{2}.$
Letting $n\to\infty$ and using Assumption 1.2, we obtain
$\mathbb{E}[Y]=\lim_{n\to\infty}\mathbb{E}[Y_{n}]\leq G^{2}\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty.$
Thus $Y$ is a nonnegative random variable with finite expectation. If $\mathbb{P}(Y=\infty)>0$, we would have $\mathbb{E}[Y]=\infty$, a contradiction. Hence
$\mathbb{P}\!\left(\sum_{t=1}^{\infty}\eta_{t}^{2}\|g_{t}\|^{2}<\infty\right)=1,$
which completes the proof. ∎
Lemma A.2.
Let $\widetilde{\Sigma}_{\infty}^{2}>1$ and define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\}.$
Fix any $L_{\alpha}\geq 1$. Then, for all $v$ with $L_{\alpha}<v\leq\widetilde{\Sigma}_{\infty}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;c_{1}\,\log\Big(\log(v+e)\Big).$
Proof.
First note that for $v>L_{\alpha}/e$ we have $v/L_{\alpha}>1/e$, hence $\log(v/L_{\alpha})>-1$, so $\log(v/L_{\alpha})+1>0$ and $\log(\log(v/L_{\alpha})+1)$ is well-defined. Moreover, $v+e>e$ implies $\log(v+e)>1$ and thus $\log(\log(v+e))>0$. We now prove the inequality. Since $L_{\alpha}\geq 1$ and the logarithm is an increasing function,
$L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty}^{2}\ \implies\ \log(v/L_{\alpha})+1\leq\log\widetilde{\Sigma}_{\infty}^{2}+1\ \implies\ \log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty}^{2}+1).$
Similarly, since $v\geq L_{\alpha}\geq 1$, we have
$v\geq 1\ \implies\ \log(v+e)\geq\log(e+1)>1\ \implies\ \log(\log(v+e))\geq\log(\log(e+1)).$
Now, by the definition of $c_{1}$,
$c_{1}\;\geq\;\frac{\log(\log\widetilde{\Sigma}_{\infty}^{2}+1)}{\log(\log(e+1))}\ \implies\ \log(\log\widetilde{\Sigma}_{\infty}^{2}+1)\;\leq\;c_{1}\,\log(\log(e+1)).$
Combining the last three displays, we obtain for all $L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty}^{2}+1)\;\leq\;c_{1}\,\log(\log(e+1))\;\leq\;c_{1}\,\log(\log(v+e)),$
which is the desired inequality. ∎
Lemma A.3.
Fix any $L_{\alpha}\geq 1$ and define $\nu_{0}:=2(L_{\alpha}+1)$. Then, for all $v\geq\nu_{0}$,
$v>L_{\alpha}+\log\Big(\log(e+v)\Big).$
Proof.
Since $L_{\alpha}\geq 1$, we have $\nu_{0}=2(L_{\alpha}+1)\geq 4$. For any $v\geq 4$, we have $e^{v}\geq e+v$, which implies that $\log(e+v)\leq v$. Since $\log(e+v)>1$, and therefore $\log\Big(\log(e+v)\Big)\leq\log v$ for all $v\geq 4$. Thus, for $v\geq\nu_{0}$, we have
$v-L_{\alpha}-\log\Big(\log(e+v)\Big)\geq v-L_{\alpha}-\log v=:g(v)$
Since $g^{\prime}(v)=1-\tfrac{1}{v}\geq 0$ for all $v\geq 1$, we have that $g$ is increasing and
$g(v)\geq g(\nu_{0})=L_{\alpha}+2-\log\Big(2(L_{\alpha}+1)\Big).$
Since $\log(2x)\leq x$ for all $x\geq 1$, we obtain
$g(\nu_{0})\geq L_{\alpha}+2-(L_{\alpha}+1)=1>0.$
Therefore $v-L_{\alpha}-\log(\log(e+v))>0$ for all $v\geq\nu_{0}$, which concludes the proof. ∎
Lemma A.4.
Let $\gamma\in(0,1)$ and define
$S_{t}^{(\gamma)}:=\sum_{s=1}^{t}s^{-\gamma}\qquad\forall t\geq 1.$
Then, for all $t\geq 1$,
$S_{t}^{(\gamma)}\;\geq\;\frac{t^{1-\gamma}-1}{1-\gamma}.$
Proof.
The function $x\mapsto x^{-\gamma}$ is positive and strictly decreasing on $1,\infty)$, for any $\gamma>0$. Therefore, for each integer $t\geq 1$,
$\sum_{s=1}^{t}s^{-\gamma}\;\geq\;\int_{1}^{t+1}x^{-\gamma}\,dx.$
Evaluating the integral for $\gamma\in(0,1)$ gives
$\int_{1}^{t+1}x^{-\gamma}\,dx\;=\;\left[\frac{x^{1-\gamma}}{1-\gamma}\right]_{x=1}^{x=t+1}\;=\;\frac{(t+1)^{1-\gamma}-1}{1-\gamma}.$
To conclude, note that $(t+1)^{1-\gamma}\geq t^{1-\gamma}$ for all $t\geq 1$ and $\gamma\in(0,1)$. ∎
Lemma A.5.
Define
$H_{t}:=\sum_{s=1}^{t}\frac{1}{s},\qquad\forall t\geq 1.$
Then for all $t\geq 1$,
$H_{t}\;\geq\;\log(t).$
Proof.
The function $x\mapsto 1/x$ is positive and strictly decreasing on $1,\infty)$. Hence, for each integer $t\geq 1$,
$\sum_{s=1}^{t}\frac{1}{s}\;\geq\;\int_{1}^{t+1}\frac{dx}{x}\;=\;\log(t+1),$
The final statement follows from the monotonicity of the logarithm: $\log(t+1)\geq\log t$ for all $t\geq 1$. ∎